# Beyond Conventional Transformers: The Medical X-ray Attention (MXA) Block for Improved Multi-Label Diagnosis Using Knowledge Distillation

### v1

@@ 7.2 Ablation Study
To isolate the effects of each component, we conduct an ablation study whose summary is shown in Table 4. We compare three different new configurations in test runs over 50 training epochs each with improved augmentations:
  1. 1.
    U1 + Better Augs: A multi-label EfficientViT trained with the U-1 label-handling scheme in CheXpert but without MXA or KD.
  2. 2.
    U1 + Better Augs + MXA: We incorporate the Medical X-ray Attention (MXA) block in parallel with the standard transformer attention to capture localized anomalies.
  3. 3.
    U1 + Better Augs + MXA + KD: We add knowledge distillation (KD) from a DenseNet-121 teacher, aiming to capture more subtle co-occurrences of pathologies.
Table 4: Ablation Study Comparison of Models Overall Across 50 Epochs
Augmentation Total Loss Micro AUC
U1 + Better Augs 0.6819 0.6659
U1 + Better Augs + MXA 0.3729 0.8344
U1 + Better Augs + MXA + KD 0.3685 0.8393
@@ 7.3 Per-Pathology Analysis
Overall, the per-pathology results confirm that both global context and local enhancement matter for chest X-ray classification. Our final approach (U1 + MXA + KD) achieves the highest AUC on the majority of pathologies, especially for those that had been historically underrepresented or more difficult to detect. In practice, we find that integrating MXA yields more interpretable attention maps, often localizing subtle regions (e.g., small lesions) that the baseline missed. Combined with knowledge distillation, the model better distinguishes confounding signs such as cardiomegaly vs. consolidated lung fields, leading to fewer false positives in real-world usage scenarios.
@@ 8 Discussion
This paper demonstrates the efficacy of augmenting general-purpose attention mechanisms with task-specific modules tailored to the unique challenges of a given domain. By introducing the Medical X-ray Attention (MXA) block and leveraging knowledge distillation, we address critical limitations in training models for medical imaging classification, such as the need for computational efficiency, subtle feature extraction, and multi-label classification. The integration of these techniques into the Efficient Vision Transformer (EfficientViT) architecture highlights the broader potential of task-specific attention mechanisms and resource-efficient training strategies for advancing AI across diverse domains.
8.1 Key Findings
Our results demonstrate that the proposed MXA block and KD framework substantially enhance the performance of EfficientViTs for multi-label chest X-ray classification. Specifically, the integration of the MXA block improves the model’s ability to capture localized abnormalities while maintaining global contextual awareness, which is evident from the significant increase in ROC AUC, rising from 0.66 in the baseline model to 0.85 in the proposed method. Furthermore, the addition of KD refines the model’s predictions, particularly for challenging pathologies. The ablation study highlights the individual contributions of the MXA and KD, with MXA alone reducing the total loss by nearly half and improving AUC from 0.666 to 0.834. Underscoring its ability to enhance feature extraction, while KD marginally boosts overall performance. It generates a greater delta in AUC among pathologies with a myriad of ambiguous or uncertain labels. Together, these components create a robust framework for multi-label classification in medical imaging.
8.2 Findings in Context
The improvements achieved by our proposed methodology are particularly significant when contextualized within the clinical standards for diagnostic models. AUC values above 0.80 are considered clinically useful [40], and our model surpasses this threshold on a micro-averaged basis. The best example is the AUC for Pleural Effusion which reached 0.91, demonstrating performance superior to some radiologists [41].
Additionally, the computational efficiency of EfficientViTs, combined with the lightweight MXA block, makes inference speed and cost suitable for real-time or resource-constrained applications, such as deployment in regions with high demand but limited healthcare infrastructure.
Lastly, compared to prior works such as CheXNet [9] and DenseNet-based models, our approach offers a more scalable solution by reducing computational overhead while maintaining high performance. The use of knowledge distillation further addresses the challenges of limited labeled data by leveraging pre-trained teacher models. Allowing our framework to effectively transfer knowledge from larger networks to more efficient architectures, enabling broader accessibility and deployment.
@@ 8.3 Limitations
Furthermore, our study was constrained by limited computational resources, which introduced additional challenges. Specifically, we were unable to generate a custom teacher model, a step that we hypothesize could have significantly enhanced the overall performance of KD. The lack of resources also restricted our ability to conduct larger-scale experiments, thereby limiting our capacity to explore the full potential of our proposed methodology. In particular, questions remain regarding how our methodology would perform under systematically optimized hyperparameters and with extended training epochs. Addressing these limitations in future work could provide deeper insights into the scalability and robustness of both our proposed methodology and the future of task-specific attention mechanisms.
@@ 8.4 Future Directions
8.5 Broder Implications
Task-specific attention mechanisms and knowledge distillation present a promising avenue for advancing AI in specialized domains. By enhancing model efficiency and performance while addressing challenges such as limited data and resource constraints, these techniques hold significant potential to improve a wide range of applications. Their adaptability to domain-specific requirements underscores their broader applicability across diverse fields, including scientific research and industry. This paper highlights the positive impact of task-specific attention mechanisms, emphasizing the need for further exploration and refinement to optimize their engineering into real-world systems.
@@ 9 Conclusion
In conclusion, this paper presents a novel framework for multi-label chest X-ray classification that combines the Medical X-ray Attention (MXA) block with knowledge distillation. By addressing the unique challenges of medical imaging, such as subtle feature extraction and label uncertainty, our approach achieves significant improvements in performance, efficiency, and feature extraction. While limitations remain, the proposed method lays the groundwork for future advancements in task-specific attention mechanisms and resource-efficient AI models. With further refinement and adaptation, this scheme has the potential to transform classification workflows across diverse domains.

### v2

@@ 5.2 Ablation Study
To isolate the contribution of each module we train three variants, each for 50 epochs with identical, strengthened augmentations: (i) U1 is the multi-label M5 EfficientViT that follows the U-1 label protocol but omits MXA and KD, (ii) U1+MXA adds the Medical X-ray Attention block in parallel with the transformer attention, and (iii) U1+MXA+KD further incorporates KD from a frozen DenseNet-121 teacher.
@@
Table 2: Ablation Study Comparison of M5 Models Overall Across 50 Epochs
Augmentation AUC
U1 + Better Augs 0.6659
U1 + Better Augs + MXA 0.8344
U1 + Better Augs + MXA + KD 0.8393
(a) Training AUC curves
(b) Final-epoch AUC bars
Figure 3: (a) Training AUC over 30 epochs. (b) Final-epoch AUC for each ablation. MXA alone delivers the largest jump in AUC; adding KD yields an additional boost, lifting performance from $0.66\rightarrow 0.85$
@@ 5.3 Per-Pathology Analysis
Overall, the per-pathology results confirm that both global context and local enhancement matter for CXR classification. Our final approach (U1 + MXA + KD) achieves the highest AUC on the majority of pathologies, especially for those that had been historically underrepresented or more difficult to detect. In practice, we find that integrating MXA yields more interpretable attention maps, often localizing subtle regions (e.g., small lesions) that the baseline missed. Combined with KD, the model better distinguishes confounding signs such as cardiomegaly vs. consolidated lung fields, leading to fewer false positives in real-world scenarios.
@@
Figure 4: After 25 training epochs, an inference pass of the improved model with the MXA yields more focused and clinically meaningful attention on a CXR with pneumonia. Each heat-map pixel is the normalized attention score for that image patch. Bright yellow in the Naive/Improved panels = high attention; dark purple = low attention. The Delta panel shows the difference in attention: red indicates regions where MXA attends less than naive MHSA, blue where it attends more, and white means no change. MXA suppresses spurious focus on the shoulders while amplifying attention over the lower left lung field where consolidation is visible, mirroring radiologist practice.
@@ 6 Discussion
This paper demonstrates the efficacy of augmenting general-purpose attention mechanisms with task-specific modules tailored to the unique challenges of a given domain. By introducing the Medical X-ray Attention (MXA) block and leveraging KD, we address critical limitations in training models for medical imaging classification, such as the need for computational efficiency, subtle feature extraction, and multi-label classification. The integration of these techniques into the Efficient Vision Transformer (EfficientViT) architecture highlights the broader potential of task-specific attention mechanisms and resource-efficient training strategies for advancing AI across diverse domains.
@@ 6.1 Limitations
Furthermore, our study was constrained by limited computational resources, which introduced additional challenges. Specifically, we were unable to generate a custom teacher model, a step that we hypothesize could have significantly enhanced the overall performance of KD. The lack of resources also restricted our ability to conduct larger-scale experiments, limiting our capacity to explore the full potential of our proposed methodology. Specifically, questions remain regarding how our methodology would perform under systematically optimized hyperparameters and with extended training epochs. Addressing these could provide deeper insights into the scalability and robustness of both our proposed methodology and the future of task-specific attention mechanisms.
@@ 7 Conclusion
We have presented a task-aware attention framework that bridges the gap between generic vision transformers and the demands of clinical X-ray diagnosis. By embedding a lightweight Medical X-ray Attention (MXA) block alongside a calibrated DenseNet-121 teacher, our EfficientViT-MultiLabel model achieves a remarkable +0.19 absolute AUC gain on CheXpert—while preserving inference cost. These results underscore the power of domain-specific attention and resource-efficient distillation. In practice, our approach can be deployed on standard hospital workstations to accelerate and augment radiology workflows. We believe this work lays a versatile blueprint for marrying transformer expressivity with practical deployment constraints in healthcare and beyond.
Acknowledgments and Disclosure of Funding
We would like to express our gratitude to Sam Lin for their help in creating Figure 1, along with a research poster, which greatly enhanced the clarity and presentation of our work. We also thank Lambda Lab support for their guidance in overcoming issues with our cloud environment. Additionally, our research was supported with Cloud TPUs from Google’s TPU Research Cloud (TRC).
Appendix A Formal definition of the U-1 label protocol
Formally, for each pathology label
$y_{c}=\begin{cases}1,&\text{label is 1 or -1 (uncertain)},\\ 0,&\text{otherwise}.\end{cases}$
Thus, each image is paired with a multi-label target $y\in\{0,1\}^{C}$, where $C=14$ corresponds to the number of binary pathology indicators provided per CXR in our CheXpert training set.
Appendix B Formal definition of binary cross-entropy with logits (BCEWL)
$\mathcal{L}_{\text{BCEWL}}(o,y)=-\!\!\sum_{c=1}^{C}\!\bigl{[}y_{c}\log\sigma(o_{c})+(1-y_{c})\log\bigl{(}1-\sigma(o_{c})\bigr{)}\bigr{]},$
where $\sigma$ is the sigmoid.
Appendix C Formal definition of ROI
Given an input feature map $F\in\mathbb{R}^{B\times C\times H\times W}$, where $B$ is the batch size, $C$ the number of channels, and $H,W$ the height and width, our ROI-pooling mechanism selects critical regions via a region predictor, formally defined as
$\mathrm{ROI}_{i}=\bigl{[}x_{1},\,y_{1},\,x_{2},\,y_{2}\bigr{]},\qquad i=1,\ldots,B,$
where $\bigl{[}x_{1},y_{1}\bigr{]}$ and $\bigl{[}x_{2},y_{2}\bigr{]}$ are the normalized top-left and bottom-right coordinates of the ROI for the $i$-th image. These coordinates are generated by a lightweight convolutional network conditioned on the feature map. The bounding-box region predictor is trained jointly with the classification objective in an end-to-end fashion: during the forward pass, each predicted ROI is used to crop and rescale the input feature map, and the resulting ROI features flow into the rest of the network for classification. The gradients from the multi-label classification loss then back-propagate through the ROI predictor as well, optimizing both the bounding boxes and the classification parameters simultaneously. Consequently, the network learns ROIs that best localize relevant abnormalities for improved multi-pathology detection.
Appendix D Formal definition of CBAM-style attention
For a given pooled feature map $F_{\text{pooled}}\in\mathbb{R}^{C\times H\times W}$, channel attention is computed as
$M_{C}\;=\;\sigma\!\Bigl{(}W_{2}\,\delta\!\bigl{(}W_{1}\,\operatorname{GAP}(F_{\text{pooled}})\bigr{)}+W_{2}\,\delta\!\bigl{(}W_{1}\,\operatorname{GMP}(F_{\text{pooled}})\bigr{)}\Bigr{)},$
where $W_{1}\in\mathbb{R}^{C\times C/r}$ and $W_{2}\in\mathbb{R}^{C/r\times C}$ are learnable weight matrices, $\delta$ denotes the ReLU activation [34], $\sigma$ is the sigmoid function [35], and $r$ is the reduction ratio controlling the dimensionality of the MLP. The resulting attention map $M_{C}\in\mathbb{R}^{C}$ is applied element-wise to the feature map:
$F_{\text{chan}}\;=\;M_{C}\odot F_{\text{pooled}}.$
Next, spatial attention identifies the most critical regions within each channel-refined feature map [8]. Using $F_{\text{chan}}$, we pool along the channel dimension with max and average pooling to obtain
$F_{\text{spatial}}\;=\;\bigl{[}\operatorname{MaxPool}(F_{\text{chan}}),\;\operatorname{AvgPool}(F_{\text{chan}})\bigr{]},\qquad F_{\text{spatial}}\in\mathbb{R}^{2\times H\times W}.$
These descriptors are concatenated and passed through a convolutional layer to produce the spatial attention map
$M_{S}\;=\;\sigma\!\bigl{(}\operatorname{Conv2D}(F_{\text{spatial}})\bigr{)},\qquad M_{S}\in\mathbb{R}^{H\times W}.$
The spatial attention map is then applied element-wise:
$F_{\text{spat}}\;=\;M_{S}\odot F_{\text{chan}},\qquad F_{\text{CBAM}}\;=\;F_{\text{spat}}.$
Appendix E Formal definition of soft knowledge distillation term
Let $O^{\text{t}}\in\mathbb{R}^{18}$ denote the teacher logits and the student outputs logits be $O^{\mathrm{s}}$. Let the adapted teacher probabilities be represented as $p^{\mathrm{t}}=\sigma(O^{\mathrm{t}})$. Thus, our soft distillation is defined as
$\mathcal{L}_{\mathrm{KD}}^{\mathrm{soft}}\;=\;\frac{1}{C}\sum_{j=1}^{C}w_{j}\,\operatorname{BCEWithLogitsLoss}\!\bigl{(}O^{\mathrm{s}}_{j},\,p^{\mathrm{t}}_{j}\bigr{)},$
with $C=14$ labels.
Appendix F Qualitative MXA ROI examples and patient-level metadata
pt1
pt2
pt3
pt4
pt5
pt6
pt7
pt8
Figure F.1: Demonstration of the MXA block. MXA consistently highlights clinically relevant regions across eight patients, confirming its ability to localize subtle abnormalities. Each panel (pt1–pt8) shows the region of interest automatically pooled after an initial inference on chest X-rays. Red boxes mark MXA-predicted ROIs. Additional pathology metadata appear in Table F.1.
Table F.1: Metadata for the 8 MXA test figures
Patient NF ECM CM LO LL ED CON PNA ATL PTX PE PO FX SD
pt1 0 1 1 1 0 0 1 1 1 0 1 0 0 0
pt2 0 1 1 1 0 1 1 1 1 0 1 0 0 0
pt3 0 0 0 1 0 0 1 1 0 0 0 0 0 1
pt4 0 0 0 1 0 0 1 1 1 0 1 0 0 0
pt5 0 0 0 1 0 0 1 1 1 0 1 1 0 1
pt6 0 1 1 1 0 0 1 1 1 0 1 0 0 0
pt7 0 1 1 1 0 0 1 1 1 0 1 0 0 1
pt8 0 0 0 1 0 1 1 1 1 0 1 0 0 0
Appendix G Additional Qualitative MXA examples on CXRs
Figure G.1: Additional attention map visualizations after 25 training epochs are shown. Across CXRs with pneumonia and diverse acquisition geometries the MXA block consistently (i) reduces attention on bony anatomy and monitoring hardware, and (ii) reallocates it toward pleural regions containing abnormalities, corroborating the quantitative AUC gains reported in Tab. 1. Bright yellow in the Naive/Improved panels = high attention; dark purple = low attention. The Delta panel shows the difference in attention: red indicates regions where MXA attends less than naive MHSA, blue where it attends more, and white means no change.
Appendix H Multi-label EfficientViT design space
Table H.1 lists the six principal variants of our multi-label EfficientViT architectures we defined. Each model family member differs in its stage-wise width ($C_{i}$), depth ($L_{i}$), and number of heads ($H_{i}$). Specifically, we break the network into three stages, and each stage’s dimensions are adjusted to control the overall capacity and computational cost. Resource limits restricted our experiments to the M5 configuration, which offers the best parameter–FLOP trade-off within that budget. Future work will explore the remaining variants.
Table H.1: Multi-label EfficientViT architecture variants
Model $\{C_{1},C_{2},C_{3}\}$$\{L_{1},L_{2},L_{3}\}$$\{H_{1},H_{2},H_{3}\}$
EfficientViT-MultiLabel-M0 {64, 128, 192} {1, 2, 3} {4, 4, 4}
EfficientViT-MultiLabel-M1 {128, 144, 192} {1, 2, 3} {2, 3, 3}
EfficientViT-MultiLabel-M2 {128, 192, 224} {1, 2, 3} {4, 3, 2}
EfficientViT-MultiLabel-M3 {128, 240, 320} {1, 2, 3} {4, 3, 4}
EfficientViT-MultiLabel-M4 {128, 256, 384} {1, 2, 3} {4, 4, 4}
EfficientViT-MultiLabel-M5 {192, 288, 384} {1, 3, 4} {3, 3, 4}
Appendix I Additional quantitative plots
Figure I.1: Mean AUC over three validation runs for baseline and proposed models; error bars denote 95% confidence intervals (±$\pm 2\sigma$). Across three runs, MXA + KD improves mean AUC drastically over the baseline, demonstrating robustness to random seed.
Figure I.2: Box-plot comparison of per-epoch AUC for three runs (baseline vs. proposed). Every epoch shows higher and less variable AUC for MXA + KD, indicating both greater accuracy and consistency than the baseline.
Appendix J Reported compute resources
Table J.1: Compute Resources and Usage
Resource Specification
Lambda Labs Node
   GPU 1$\times$ NVIDIA H100 (80 GB PCIe)
   CPU 26 vCPUs
   Memory 200 GiB RAM
   Storage 1 TiB SSD
   Number of experiments $\sim$120
   Avg. compute time per experiment $\sim$10 hours
   Total compute time $\sim$1,200 hours
Google TPU
   TPU 1$\times$ v4-8
   Number of experiments $\sim$1
   Avg. compute time per experiment $\sim$50 hours
   Total compute time $\sim$ 50 hours